# Neil A. Chriss

### v2

@@ 6.6 What is really going on with strategic centralization?
The conclusion of the prior sections on centralization is that a firm can improve its implementation costs, but it requires an artful masking of the true number of traders to optimally minimize costs. The question is what is really going on and how is this possible at all?
@@
To start, note that as a system the total cost of trading varies depending on the number of traders. So when a firm masks its true number of traders it is in effect changing the game that is being played itself. Ordinarily this is not possible because each trader is totally independent of all the others. However, when there is a firm involved and that firm employs a number of traders then the firm can effectively modify the number of traders in the market to its advantage. Exactly what that advantage is depends on two opposing forces.
The first of those forces is that as the number of traders in competition increases, all else equal the total cost of trading increases as well. This is a consequence of the degradation of efficiency a system undergoes in a non-cooperative game. The second is that for a given trader, their share of the total cost changes when the number of traders changes. When the number of traders is very small (two or three, say) then increasing the number of traders increases share of cost for a given trader and this increase is significantly greater for traders seeking a large fraction of the total quantity. On the other hand, for a moderate number of traders, the reverse of true and increasing traders decreases cost share.
@@
7 Acknowledgements
I extend my gratitude to the Machine Learning Research Group at Morgan Stanley and Vlad Ragulin who provided much discussion and feedback. I would also like to thank Michael Kearns and Mirah Shih of the University of Pennsylvania for extremely valuable discussions concerning algorithmic game theory.
@@ 7 Acknowledgements
To prove the equality in Eq. 4.7a we use the Euler-Lagrange equation on the loss function $C_{i,\lambda_{i}}=(\dot{m}+\kappa m)\lambda_{i}a_{i}$. First we expand the loss function:
$C_{i,\lambda_{i}}=\left(\sum_{j}\lambda_{j}\dot{a}_{j}+\kappa\sum\lambda_{j}a_{j}\right)\lambda_{i}\dot{a}_{i}$ (A.1)
@@
$\begin{split}\frac{\partial}{\partial t}\frac{\partial C_{i,\lambda_{i}}}{\partial\dot{a}_{i}}&=\sum_{j\neq i}^{n}\lambda_{i}\lambda_{j}\ddot{a}_{j}+2\lambda_{i}^{2}\ddot{a}_{i}+\kappa\sum_{j=1}^{n}\lambda_{i}\lambda_{j}\dot{a}_{j}\\[5.0pt] \frac{\partial C_{i,\lambda_{i}}}{\partial a_{i}}&=\kappa\lambda^{2}_{i}\dot{a}_{i}\end{split}$ (A.2)
The Euler-Lagrange equations says that $\frac{\partial}{\partial t}\frac{\partial C_{i,\lambda_{i}}}{\partial\dot{a}_{i}}-\frac{\partial C_{i,\lambda_{i}}}{\partial a_{i}}=0$ which implies:
@@ A.4 Proof of equilibrium cost of trading
We immediately see that since $\sum\lambda_{i}=1$, that $\sum(\lambda_{i}n-1)=0$ and thus

### v3

@@ 6.6 What is really going on with strategic centralization?
The conclusion of the prior sections is that a firm can reduce its implementation costs by centralizing and matching the number of non-firm traders with the number of traders it represents to the market. This is the content of Eq. 6.14 and Eq. 6.15. As already noted it is surprising that this value is independent of both the firm’s total target quantity and the alpha-decay parameter. However, our question here is, what is really going on and how is this possible at all? First, we remind the reader that there is an optimal number of traders for a firm to represent when centralizing and this is given by (6.14). In this section we discuss why this is the case.
@@
To start, note that as a system the total cost of trading varies depending on the number of traders. So when a firm masks its true number of traders it is in effect changing the game that is being played itself. Ordinarily this is not possible because each trader is totally independent of all the others. However, when there is a firm involved and that firm employs a number of traders then the firm can effectively modify the number of traders in the market to its advantage.
Exactly what that advantage is depends on two opposing forces. The first of those forces is that as the number of traders in competition increases, all else equal the total cost of trading increases as well. This is a consequence of the degradation of efficiency a system undergoes in a non-cooperative game. The second is that for a given trader, their share of the total cost changes when the number of traders changes. When the number of traders is very small (two or three, say) then increasing the number of traders increases share of cost for a given trader and this increase is significantly greater for traders seeking a large fraction of the total quantity. On the other hand, for a moderate number of traders, the reverse of true and increasing traders decreases cost share.
To see how this plays out, first observe that as the number of traders competing for a stock increases, the total cost over all traders increases. This is another example of the general efficiency degradation in systems where individuals act in their own interest. Figure 13 shows the percentage change in total cost of trading versus a baseline of two traders as the number of traders overall grows.
Figure 13: The relationship between the total number of traders competing for a stock and the percentage increase in the total cost of trading (over all traders) versus when there are two traders. We see that in all cases the shape of the increase is the same and is quite sharp till approximately 20 traders. The difference between each plot, the level of alpha-decay has a significant impact on the level of the increase. We see that in the case of very high alpha decay ($\kappa=25$) costs rise to almost 50% greater than base case.
Given Figure 13 we begin to see why there is a potential benefit to strategic centralization. Given a starting situation, where for example a firm begins 25 traders and there are another 10 outside of the firm, then centralization allows a firm to in effect change where the entire competition is on the graph. For example if the firm consolidates from 25 traders to a single trader (as in naive centralization) then the firm can effectively change the total number of traders from 35 to 11, sharply decreasing the total number of traders and moving the total cost down substantially. How much impact this has overall depends on the level of $\kappa$. If this were the only force in play then naively centralizing all trades would indeed lower costs as the total cost over all traders decreases. However, the second force at play increases the relative share of total costs to a given trader as that trader becomes a greater fraction of all traders. This is the message of Figure 5. Thus there is an optimal balance to be achieved. One can see this in Figure 14.
@@
7 Acknowledgments
I extend my gratitude to the Machine Learning Research Group at Morgan Stanley, especially Yuriy Nevmyvaka for some very useful observations, and Vlad Ragulin who provided much discussion and feedback. I would also like to thank Michael Kearns and Mirah Shih of the University of Pennsylvania for extremely valuable discussions concerning algorithmic game theory.
@@ 7 Acknowledgments
To prove the equality in Eq. 4.7a we use the Euler-Lagrange equation on the loss function $C_{i}=(\dot{m}+\kappa m)\lambda_{i}a_{i}$. First we expand the loss function:
$C_{i}=\left(\sum_{j}\lambda_{j}\dot{a}_{j}+\kappa\sum\lambda_{j}a_{j}\right)\lambda_{i}\dot{a}_{i}$ (A.1)
@@
$\begin{split}\frac{\partial}{\partial t}\frac{\partial C_{i}}{\partial\dot{a}_{i}}&=\sum_{j\neq i}^{n}\lambda_{i}\lambda_{j}\ddot{a}_{j}+2\lambda_{i}^{2}\ddot{a}_{i}+\kappa\sum_{j=1}^{n}\lambda_{i}\lambda_{j}\dot{a}_{j}\\[5.0pt] \frac{\partial C_{i}}{\partial a_{i}}&=\kappa\lambda^{2}_{i}\dot{a}_{i}\end{split}$ (A.2)
The Euler-Lagrange equations says that $\frac{\partial}{\partial t}\frac{\partial C_{i}}{\partial\dot{a}_{i}}-\frac{\partial C_{i}}{\partial a_{i}}=0$ which implies:
@@ A.4 Proof of equilibrium cost of trading
Note that because $\sum_{i=1}^{n}\lambda_{i}=1$, we have $\sum_{i=1}^{n}(\lambda_{i}n-1)=n\sum_{i=1}^{n}\lambda_{i}-\sum_{i=1}^{n}1=0$. Therefore,
@@ A.5 Implementation cost share
Appendix B Proof of minimizing implementation costs via centralization
We wish to minimize the expression in Eq. 6.13 with respect to $m$. We treat $m$ as a continuous variable. To find the minimum of the expression with respect to $m$:
$E(m)=\kappa\frac{\lambda_{n_{1}}(n+m)-(n_{1}+m)}{(n+m)(1-e^{-\kappa})}+\kappa\frac{n_{1}+m}{(n+m)(e^{\kappa}-1)}+\frac{(n_{1}+m)\kappa}{n+m+1}$
we can re-parameterize it by setting $s=n+m$, so $m=s-n$. This transformation allows us to rewrite the expression in terms of $s$:
$E(s)=\kappa\left(\frac{\lambda_{n_{1}}s-(n_{1}+s-n)}{s(1-e^{-\kappa})}+\frac{n_{1}+s-n}{s(e^{\kappa}-1)}+\frac{n_{1}+s-n}{s+1}\right).$
To simplify further, let $n_{2}=n-n_{1}$, then $n_{1}+s-n=s-n_{2}$, and the expression becomes:
$E(s)=\kappa\left(\frac{(\lambda_{n_{1}}-1)s+n_{2}}{s(1-e^{-\kappa})}+\frac{s-n_{2}}{s(e^{\kappa}-1)}+\frac{s-n_{2}}{s+1}\right)$ (B.1)
We now take the derivative of $E(s)$ with respect to $s$ by differentiating each term in the sum individually. Writing $E_{1},E_{2}$ and $E_{3}$ for the three summands in Eq. B.1 we have:
$C_{1}(s)$ $=\kappa\frac{(\lambda_{n_{1}}-1)s+n_{2}}{s(1-e^{-\kappa})}\quad\implies\quad C_{1}^{\prime}(s)=-\kappa\frac{n_{2}}{s^{2}(1-e^{-\kappa})}$ (B.2a)
$C_{2}(s)$ $=\kappa\frac{s-n_{2}}{s(e^{\kappa}-1)}\quad\implies\quad C_{2}^{\prime}(s)=\kappa\frac{n_{2}}{s^{2}(e^{\kappa}-1)}$ (B.2b)
$C_{3}(s)$ $=\kappa\frac{s-n_{2}}{s+1}\quad\implies\quad C_{3}^{\prime}(s)=\kappa\frac{1+n_{2}}{(s+1)^{2}}$ (B.2c)
We can see clearly that non of the three summands varies with respect to $s$ and this explains why the cost minima do not depend on number of traders added (where $m$ is number of traders added and $s=m+n$). Combining these results, we get:
$C^{\prime}(s)=C_{1}^{\prime}(s)+C_{2}^{\prime}(s)+C_{3}^{\prime}(s)=\kappa\left(-\frac{n_{2}}{s^{2}(1-e^{-\kappa})}+\frac{n_{2}}{s^{2}(e^{\kappa}-1)}+\frac{1+n_{2}}{(s+1)^{2}}\right)$ (B.3)
We simplify the first two terms in $C^{\prime}(s)$ using the identity $e^{\kappa}-1=(1-e^{-\kappa})e^{\kappa}$:
$-\frac{n_{2}}{s^{2}(1-e^{-\kappa})}+\frac{n_{2}}{s^{2}(e^{\kappa}-1)}$ $=-\frac{n_{2}e^{\kappa}}{s^{2}(e^{\kappa}-1)}+\frac{n_{2}}{s^{2}(e^{\kappa}-1)}$
$=-\frac{n_{2}e^{\kappa}-n_{2}}{s^{2}(e^{\kappa}-1)}$
$=-\frac{n_{2}}{s^{2}}$
Thus, the derivative simplifies to:
$C^{\prime}(s)=\kappa\left(-\frac{n_{2}}{s^{2}}+\frac{1+n_{2}}{(s+1)^{2}}\right).$ (B.4)
We now set $C^{\prime}(s)=0$ and obtain:
$-\frac{n_{2}}{s^{2}}+\frac{1+n_{2}}{(s+1)^{2}}=0.$ (B.5)
Multiplying both sides by $s^{2}(s+1)^{2}$ to clear the denominators, we obtain:
$-n_{2}(s+1)^{2}+(1+n_{2})s^{2}=0.$ (B.6)
which in turn implies:
$-n_{2}(s^{2}+2s+1)+(1+n_{2})s^{2}=0,$ (B.7)
and this simplifies to:
$s^{2}-2n_{2}s-n_{2}=0$ (B.8)
We may solve it using the quadratic formula:
$s=\frac{2n_{2}\pm\sqrt{(2n_{2})^{2}+4n_{2}}}{2}=n_{2}+\sqrt{n_{2}^{2}+n_{2}}$ (B.9)
Since $s=n+m$ (the original number of traders modified by $m$), it must be positive, and therefore we take the positive root:
$s=n_{2}+\sqrt{n_{2}^{2}+n_{2}}$ (B.10)
Finally since $s=n+m$ we have Recalling that $m=s-n$ and $n_{2}=n-n_{1}$, we have:
$m=s-n=-n_{1}+\sqrt{(n-n_{1})(n-n_{1}+1)}.$ (B.11)
and we conclude that the minimum of $E(m)$ occurs at:
$m=-n_{1}+\sqrt{n_{2}(n_{2}+1)}$ (B.12)
and this proves the result. $\blacksquare$